If the least squares equation for a company time series starting in 2011 and ending in...

  1. If the least squares equation for a company time series starting in 2011 and ending in 2018 is Ŷ = 12.5 + 1.87t (in millions), what is the t and estimate for 2023?
Answer #1

The estimate starts with 2011 so t = 0 for 2011. Hence,

For 2023, t = 2023 - 2011 = 12

Also,

Estimate for 2023 = 12.5 + 1.87*12 = 34.94 millions